# “Did They Consent to That?”: Safer Digital Intimacy via Proactive Protection Against Image-Based Sexual Abuse

### v1

@@ 5 Sharing Contexts
Established relationship. A vast majority of participants shared intimate content with those they had established relationships (e.g. significant other, spouse). In line with literature [48, 17], participants frequently mentioned the importance of intimate content as a form of connecting with their partners, “intimate content is to keep a connection alive, it’s part of a couple’s sex life” (P48). Around a quarter of participants also shared content with friends as a form of bonding. “I have been in text groups with like, female friends, and we’ve sent photos of our bodies to each other, just for fun, or like, hey, I’m feeling sexy today, or look at this great photo I took” (P18).
@@
Strangers. Individuals may also share intimate content with those they do not know and with whom they are not intending to initiate a relationship. For example, participants shared content mutually with strangers (e.g., in online communities) or in a one directional sharing relationship, such as with a social media audience. Our recreational participants were motivated by a desire to connect with others, artistic expression, or to receive body validation or express body positivity [48, 81]. For example, P26 shared that, “Early [in] my transition, I had the need to just share some pictures [to] feel more sexy…I created an account [on a social media platform] especially for that, that [account] is networked with mostly other trans women that also have accounts for the exact same purpose.”
@@ 6.1 Recipient Resharing
Despite many participants expressing that they trusted their recipients, they reported having passive or background fears about their content being reshared. P31 described , “[in] that moment before [I] fall asleep, [I feel] like 'Oh, shit,' all of this is out there. [I] have a bit of a panic. But most of the time, I would say I feel relatively confident that if it does come out, I [will] deal with it in a way that I find suitable.” Particularly for those who had experienced NDII previously, recipient resharing was a primary concern. A participant explained: “I don’t think there’s ever been a time where I share content and I don’t think about the harm that can be caused” (P35). Participants also reported secondary concerns such as not knowing they had been victimized: “I know there are groups on Telegram and WhatsApp where people send porn. And then you don’t have any control over it. Because if you’re not part of these groups, you don’t know they’re there” (P9). In addition, P3 noted the concern of aggregator pages further proliferating nonconsensual content by reposting it. News articles and research on the existence of anonymous forums and messaging groups that are dedicated to noncensually sharing intimate content support these fears [71, 38, 9]. Participants also spoke about fears of an ex-partner sharing intimate content after a breakup.“I asked him to delete all my intimate images and videos, but I don’t know if he really did because we had a hard break up” (P14).
@@ 6.2 Non-Recipient Threats
Data Breaches & Hacking. Although many described hacking as “a hypothetical background fear” (P11), more than a third were concerned about it as a means of accessing their stored intimate content. They brought up risks on their personal devices—of viruses, phishing scams, and malware—and the risk of their intimate content being leaked through a data breach or poor security practices/policies of a platform they had used for sharing (e.g. how content is stored, protected, deleted). People feared both technical and policy breaches; exemplifying the latter, P10 “read stories of [employees] who have set up programs to find [images] of women shirtless.”
@@ 6.2.1 Unintentional Discovery
Device Theft or Loss. More than a sixth feared that someone could access stored intimate content if a device were stolen or lost. This was often coupled with concerns of weak password security: “someone like could crack my code and take my pictures” (P30). P51 experienced such an incident and described, “once my phone was stolen, containing a lot, a lot, a lot of personal images and data… I was like, 'Oh, my [gosh]. Oh, my [gosh]'.” Additionally, P38 asked a recipient to delete any intimate content prior to bringing a phone in for repair—a concern confirmed by cases alleging intimate content theft by customer service and technical support employees who were given device access [27, 28, 16].
@@ 6.3 Mechanism of Harm: Stigma
IBSA can lead to many forms of harm (see Section 2). that stigma and victim blaming can perpetuate. Due to stigma, participants feared awkward or even dangerous situations with family, friends, their communities, places of work and study. P18 worried about the reactions of their colleagues: “I would like to think that it would be a non issue, and the focus will be on the person who was nonconsensually sharing images of me. But unfortunately, there is a lot of victim blaming.”
Consistent with prior work [84, 67, 40, 83, 22], the concerns described tended to depend on participant identity. For LGBTQ and sex-working participants, the risk was magnified if they or their partner were not out, or their identity was a personal risk to them or their recipient(s). For example, P10 took more precautions because, “[My ex-partner] was from a quite conservative religious family… it’s always better to play safe, especially when your partner is closeted." In the context of sex work, P49 explained: “I’m afraid that if I do more content… that someone from my family would see that… It’s the most stressful thing for me,” and as P35 described, “I have another [non sex work] job and I [get] really scared that I will no longer get work… [despite my] paranoia, there’s nothing I can really do about it, because I have to still do [sex work].” Cultural norms also magnified risks: “Turkey is a very conservative and traditional country. My parents are open minded but I don’t think they could handle this” (P14).
Contrastingly, several cisgender men felt that they faced fewer threats: “as someone who presents as male, I think I’m less at risk from exposure than someone who presents as anywhere else on the gender spectrum, because I think my image is less likely to be shared around maliciously” (P2).
Some also felt that if intimate content was less stigmatized in their specific social community, they may receive more support if NDII occurred (though they may still have concerns about the response of those outside of that community). P26 commented that due to shared politics, “I trust the people at my workplace to not let that have a negative influence."
@@ 7 Technologies
Participants used six categories of sharing technologies: messaging (apps and SMS), social media, hookup/dating apps, adult content platforms, file share platforms, and email. Nearly all used multiple technologies as summarized by P12: “…I’ve got like Telegram, I’ve got Signal, I’ve got WhatsApp, I’ve got Kik, you name it, I’ve got it… if there’s something really large, that requires Dropbox… I’ll do that” (P12).
In total, participants used 40+ technologies for sharing intimate content. Table 1 categorizes platforms and their use cases. With the exception of 5 participants, nearly all stored intimate content locally in their mobile device photo gallery, laptop, or desktop computer. More than half intentionally stored their content in a cloud/file share platform. The vast majority used multiple forms of storage: “I have everything stored everywhere, I have it in my phone, I have it in my hard-drive. I have some parts on the cloud” (P42).
@@ 7.1 Convenience
Around half of our participants shared intimate content using the same messaging service that they were already using to communicate with others. Since participants often shared intimate content in the middle of a conversation, they emphasized the importance of conversation continuity: “…it’s like at the heat of the moment… so to [choose another method and] make it longer… it just doesn’t make any sense" (P43). About a quarter also commented on the importance of usability and highlighted different affordances that they valued: fast message delivery, desktop integration, clean user interfaces, and familiarity. Therefore individuals that prioritize convenience must make do with any existing safety tools available on the platforms they use for sharing intimate content.
@@ 7.3 Interpersonal
Number of Active Users. About a sixth of our participants preferred not to use less popular platforms because they did not have enough recipients on that platform, which inhibited the convenience of use. For example, P9 tried an app that they felt was safer but “I only had two people there. So I deleted it.” This can hinder adoption of platforms that participants may prefer for security reasons like offering E2EE.
Preference of Recipient. In the context of direct messaging, more than a quarter of participants were open to or deferred to (sometimes feeling like they had to) the technology preferred by their recipient. Particularly when the recipient prefers an app with specific safety features, some participants were willing to accommodate them. Sex-working participants more commonly expressed that they needed to use technologies that felt more secure and convenient to their clients and also because in some cases, “they’re not willing to like to download like another app or like check another website” (P32).
@@ 7.4 Commercial
Profitability. A third of sex-working participants emphasized the profit potential of a technology such as a platform’s ability to grow their client base, flexibility in setting rates (P20) and/or easy processes for receiving and transferring payments (P9, P24). Exemplifying the strong influence of profitability factors, P42 remarked, “I think [adult content platform] is the worst in everything… But I use it a lot because it’s the one that has the most people.”
@@ 8 Safety Maze: How People Try to Defend
All participants adopted at least one strategy but many relied on multiple. Since many prioritized convenience and the preference of a recipient, most participants leverage any technological features available in the sharing platform and fill in significant gaps in availability of desired features (see Section 9.1) with interpersonal behavioral strategies and third-party tools. In this section, we describe participant strategies based on the threats they sought to defend against.
8.1 Defending Against Recipient Resharing.
To prevent recipients from resharing intimate content, participants employed strategies across the timeline of first establishing trust with a recipient to after content had been shared.
@@ 8.1.1 Before Sharing
Rule Setting. About half of participants explicitly asked recipients not to share or store intimate content. “…I asked them not to screenshot anything, just to [keep] it in chat. And that was kind of the ground rule” (P11). Sex workers were more likely to set explicit rules. P35 shared: “before I send images, I always just say that this image is between myself and yourself… No one else is paying for this except you.” Recreational sexters mentioned that they would circumstantially ask for content not to be shared such as if they had friends in common (P4) or if there were identifying features in the content (P14).
In addition, a few participants stated direct consequences if content were to be shared. For instance, P30 said, “I would just tell them, if you ever shared, I will come up to you, we will go to the police” whereas P34 sent reminders to clients about potential consequences: “I send them the penal law articles regarding sharing/publishing porn content.”
@@ 8.1.2 While Sending
When available, a few participants also made use of screenshot prevention in which screenshots result in a blank screen or deterrence messaging from a platform. P15 described using a platform because, “you couldn’t also take like screenshot, like it wasn’t allowed like like how it is [on a banking app]." Since this feature was not readily available in most platforms used for sharing content (see Appendix B), it was also commonly mentioned by others as a feature they wished more platforms would implement it as a means of increasing friction, despite many noting that it could be bypassed (e.g. capturing content through another device).
@@ 8.1.3 After Sharing
Participants proactively used strategies to prevent being identified in the event of NDII. This may be to create plausible deniability (e.g. against blackmail) and to protect against additional harm (e.g. doxxing, harassment). In many cases, participants employed these strategies with the understanding that once content was shared, it was out of their control. “I know that I don’t have full control over where these images will go, if people can trace them back to me" (P26).
@@
Metadata Removal. Six participants used metadata removal to eliminate extra information that is added when an image is created (e.g. location, date/time of creation). However, they generally expressed that they would prefer for it to be automatically done by a platform: “[it] almost feels like that should be the default when you share any image anyway” (P6).
8.3 Defending Against Non-Recipients.
@@
Consent. Many participants mentioned the importance of consent prior to sharing intimate content. A few specifically mentioned in-the-moment consent checks to check that a recipient is alone to prevent shoulder surfing, “I’ll always ask first and wait for their response before I send something because you never know what people are doing” (P29).
Masking Content. Another strategy a few participants used to protect against shoulder surfers was to mask it upon sending it. For the two participants who used this feature, a recipient would receive a blurred image that they would then have to click on to reveal the contents. This would signal that the content was sensitive and should be viewed in a private location. Though this feature was not commonly mentioned, it is also not readily available in most platforms (see Section 9.1).
End-to-End Encryption. End-to-end encryption (E2EE) ensures that only a sender and their intended recipient(s) can view a message. Once an E2EE message has been delivered, it carries similar risks to unencrypted messages: the recipient can still store or share it (some platforms may auto-download images onto devices). Over a quarter of participants actively used E2EE or considered it to be useful (excluding those who happened to use platforms with E2EE but did not mention it). For a few, E2EE alleviated their concerns about unwanted access by a platform or government entity: “There are basically two possible attackers. Like either the person I’m sending this to or someone from the outside trying to break into systems… I mostly care about someone from the outside able to steal pictures” (P26). P10 added that they valued E2EE because, “if there’s anything I don’t want, it’s the government knowing details of my personal life and seeing my tits.” On the other hand, a sixth of participants did not trust E2EE, believing that it could be broken if needed: “If someone wants the data, they will get their hands on it anyway. And if that someone is the police, they will just force companies to give that to them. No matter how well encrypted it is” (P50).
@@ 8.4 Comparing Participant Groups
We encountered narratives in which experiencing harm had increased fears of NDII: that, "I’m just really paranoid… and there’s nothing I can really do about it… Especially after it’s already happened" (P35). We also observed that some victim-survivors implemented additional strategies after experiencing harm. As P21 explained, “What happened to me was forever… And so I was more knowledgeable about what was going on, about the privacy of these pictures.” P43 shared, "future photos… were really less revealing than they used to… you cannot recognise me." Otherwise, we did not observe notable differences in practices. Since we interviewed victim-survivors who continued sharing intimate content, many expressed that they resolved not to let shame and stigma stop them (and for some sex workers, they were unable to due to financial reasons). For example, P18 shared, "I’m having fun. I’m an adult… I don’t want to allow shame or stigma to prevent me from living my life and exploring my sexuality."
When comparing commercial and recreational participants, commercial participants shared content in all the same contexts along with specific platforms to sell content and/or advertise services. However, sex workers were more constrained by the preferences of their clients when selecting a sharing method. Commercial participants may also prefer sex work specific platforms since they may otherwise perceive or experience hostility. We also observed that sex workers were less worried about unintentional discovery, either since the people around them already knew they were sex workers, or they employed more strategies to prevent this.
Generally, we observed that while recreational participants adopted many strategies, sex-working participants adopted more (in particular, watermarking, removing metadata, and using safe storage methods). They were also more likely to set explicit boundaries around sharing or storing. This greater protective action may be due both to differences in threat models (e.g. greater personal/physical risk, stigma) as well as better advice and educational support (see Section 9.2).

### v2

@@ 5 Sharing Contexts
Established relationship. A vast majority of participants shared intimate content in the context of an established relationship (e.g., spouse, long-term client). In line with literature [48, 17], participants frequently mentioned the importance of intimate content as a form of connecting with their partners, “intimate content is to keep a connection alive, it’s part of a couple’s sex life” (P48). Around a quarter of participants also shared content with friends as a form of bonding. “I have been in text groups with like, female friends, and we’ve sent photos of our bodies to each other, just for fun, or like, hey, I’m feeling sexy today, or look at this great photo I took” (P18).
@@
Strangers. Individuals may also share intimate content with those they do not know and with whom they are not intending to initiate a relationship. For example, participants shared content mutually with strangers (e.g., in online communities) or in a one directional sharing relationship, such as with a social media audience. Our recreational participants were motivated by a desire to connect with others, express themselves artistically, receive body validation or express body positivity [48, 81]. For example, P26 shared that, “Early [in] my transition, I had the need to just share some pictures [to] feel more sexy…I created an account [on a social media platform] especially for that, that [account] is networked with mostly other trans women that also have accounts for the exact same purpose.”
@@ 6.1 Recipient Resharing
Despite many participants expressing that they trusted their recipients, they reported having passive or background fears about their content being reshared. P31 described, “[in] that moment before [I] fall asleep, [I feel] like 'Oh, shit,' all of this is out there. [I] have a bit of a panic. But most of the time, I would say I feel relatively confident that if it does come out, I [will] deal with it in a way that I find suitable.” Particularly for those who had experienced NDII previously, recipient resharing was a primary concern. A participant explained: “I don’t think there’s ever been a time where I share content and I don’t think about the harm that can be caused” (P35). Participants also reported secondary concerns such as not knowing they had been victimized: “I know there are groups on Telegram and WhatsApp where people send porn. And then you don’t have any control over it. Because if you’re not part of these groups, you don’t know they’re there” (P9). In addition, P3 noted the concern of aggregator pages further proliferating nonconsensual content by reposting it. News articles and research on the existence of anonymous forums and messaging groups that are dedicated to noncensually sharing intimate content support these fears [71, 38, 9]. Participants also spoke about fears of an ex-partner sharing intimate content after a breakup  [^6] See [coduto_delete_2024] for an in-depth analysis on how people manage intimate content after experiencing a breakup.: “I asked him to delete all my intimate images and videos, but I don’t know if he really did because we had a hard break up” (P14).
@@ 6.2 Non-Recipient Threats
Data Breaches & Hacking. Although many described hacking as “a hypothetical background fear” (P11), more than a third were concerned about it as a means of accessing their stored intimate content. They brought up risks on their personal devices (e.g. viruses, phishing scams, and malware) and the risk of their intimate content being leaked through a data breach or poor security practices/policies of a platform they had used for sharing (e.g. how content is stored, protected, deleted). People feared both technical and policy breaches; exemplifying the latter, P10 “read stories of [employees] who have set up programs to find [images] of women shirtless.”
@@ 6.2.1 Unintentional Discovery
Device Theft or Loss. More than a sixth feared that someone could access stored intimate content if a device were stolen or lost. This was often coupled with concerns of weak password security: “someone like could crack my code and take my pictures” (P30). P51 experienced such an incident and described, “once my phone was stolen, containing a lot, a lot, a lot of personal images and data… I was like, 'Oh, my [gosh]. Oh, my [gosh]'.” Additionally, P38 asked a recipient to delete any intimate content prior to bringing a phone in for repair—a concern confirmed by cases of intimate content theft by customer service and technical support employees who were given device access [27, 28, 16].
@@ 6.3 Mechanism of Harm: Stigma
IBSA can lead to many forms of harm (see Section 2) that stigma and victim blaming can perpetuate. Due to stigma, participants feared awkward or even dangerous situations with family, friends, their communities, and/or places of work and study. P18 worried about the reactions of their colleagues in the event of NDII: “I would like to think that it would be a non issue, and the focus will be on the person who was nonconsensually sharing images of me. But unfortunately, there is a lot of victim blaming.”
Consistent with prior work [84, 67, 40, 83, 22], the concerns described tended to depend on participant identity. For LGBTQ and sex-working participants, the risk was magnified if they or their partner were not out, or their identity was a personal risk to them or their recipient(s). For example, P10 took more precautions because, “[My ex-partner] was from a quite conservative religious family… it’s always better to play safe, especially when your partner is closeted." In the context of sex work, P49 explained: “I’m afraid that if I do more content… that someone from my family would see that… It’s the most stressful thing for me,” and as P35 described, “I have another [non sex work] job and I [get] really scared that I will no longer get work… [despite my] paranoia, there’s nothing I can really do about it, because I have to still do [sex work].” Cultural norms also magnified risks: “Turkey is a very conservative and traditional country. My parents are open-minded but I don’t think they could handle this” (P14).
In contrast, several cisgender men felt that they faced fewer threats [^7] We note that men are disproportionately affected by sextortion, another form of IBSA, in which individuals are targeted by fake online personas to begin a relationship. Once intimate content is exchanged, the content is used as leverage for financial extortion [83].: “as someone who presents as male, I think I’m less at risk from exposure than someone who presents as anywhere else on the gender spectrum, because I think my image is less likely to be shared around maliciously” (P2). Some also felt that if sharing intimate content were less stigmatized in their specific social community, they may receive more support if NDII occurred (though they may still have concerns about the response of those outside of that community). P26 commented that due to shared politics, they trusted “people at my workplace to not let [nonconsenusally shared intimate content] have a negative influence."
Sharing Method Sharing Use Case # of Platforms
Messaging 1-1 conversation or with groups of recipients 9
Social Media 1-1 conversations (DMs), groups (posting or DMs), for marketing (commercial) 10
Adult content platforms Posting content to platforms, livecamming 11
Dating / hookup apps Sharing in-app intimate content, posting semi-nude/lewd photos in profile 10
Cloud/file share platforms Storing and directly sharing content with recipients 4
Email Sharing larger files and/or links to files, occasionally used as an alternative to primary sharing technology N/A
@@ 7 Technologies
In total, participants mentioned using 40+ different platforms for sharing and storing intimate content. We categorized them in Table 1, which also includes the use cases people mentioned in relation to each category of sharing method and the number of different platforms that were mentioned per category. This leads us to emphasize that any platform that enables the storing and sharing of photo/video content is likely being used for intimate content. In addition, each participant commonly used multiple technologies as summarized by P12:
“…I’ve got like Telegram, I’ve got Signal, I’ve got WhatsApp, I’ve got Kik, you name it, I’ve got it… if there’s something really large, that requires Dropbox… I’ll do that”.
With the exception of 5 participants, nearly all stored intimate content locally in their mobile device photo gallery, laptop, or desktop computer. More than half intentionally stored their content in a cloud/file share platform. The vast majority used multiple forms of storage: “I have everything stored everywhere, I have it in my phone, I have it in my hard-drive. I have some parts on the cloud” (P42).
@@ 7.1 Convenience
Around half of our participants shared intimate content using the same messaging service that they were already using to communicate with others. Since participants often shared intimate content in the middle of a conversation, they emphasized the importance of conversation continuity: “…it’s like at the heat of the moment… so to [choose another method and] make it longer… it just doesn’t make any sense" (P43). About a quarter also commented on the importance of usability and highlighted different affordances that they valued such as fast message delivery, desktop integration, clean user interfaces, and familiarity. Therefore, individuals that prioritize convenience must make do with any existing safety tools available on the platforms they use for sharing intimate content.
@@ 7.3 Interpersonal
Number of Active Users. About a sixth of our participants preferred not to use less popular platforms because they did not have enough recipients on that platform, which inhibited the convenience of use. For example, P9 tried an app that they felt was safer but “I only had two people there. So I deleted it.” This can hinder adoption of platforms that participants may prefer for security reasons, like offering E2EE.
Preference of Recipient. In the context of direct messaging, more than a quarter of participants were open to or deferred to (sometimes feeling like they had to) the technology preferred by their recipient. Particularly when the recipient prefers an app with specific safety features, some participants were willing to accommodate them. Sex-working participants more commonly expressed that they needed to use technologies that felt more secure and convenient to their clients and also because in some cases, “they’re not willing [to] download like another app or like check another website” (P32).
@@ 7.4 Commercial
Profitability. A third of sex-working participants emphasized the profit potential of a technology, such as a platform’s ability to grow their client base, flexibility in setting rates (P20) and/or easy processes for receiving and transferring payments (P9, P24). Exemplifying the strong influence of profitability factors, P42 remarked, “I think [adult content platform] is the worst in everything… But I use it a lot because it’s the one that has the most people.”
@@ 8 Safety Maze: How People Try to Defend
All participants adopted at least one strategy but many relied on multiple. Since many prioritized convenience and the preference of a recipient, most participants leveraged any technological features available in the sharing platform and fill in significant gaps in availability of desired features (see Section 9.1) with interpersonal behavioral strategies and third-party tools. In this section, we describe participant strategies based on the threats they sought to defend against.
8.1 Defending Against Recipient Resharing
To prevent recipients from resharing intimate content, participants employed strategies across the timeline from first establishing trust with a recipient to after content had been shared.
@@ 8.1.1 Before Sharing
Rule Setting. About half of participants explicitly asked recipients not to share or store intimate content. “…I asked them not to screenshot anything, just to [keep] it in chat. And that was kind of the ground rule” (P11). Sex workers were more likely to set explicit rules. P35 shared: “before I send images, I always just say that this image is between myself and yourself… No one else is paying for this except you.” Recreational sexters mentioned that they would circumstantially ask for content not to be shared such as if they had friends in common (P4) or if there were identifying features in the content (P14). In addition, a few participants stated direct consequences if content were to be shared. For instance, P30 said, “I would just tell them, if you ever shared, I will come up to you, we will go to the police” whereas P34 sent reminders to clients about potential consequences: “I send them the penal law articles regarding sharing/publishing porn content.”
@@ 8.1.2 While Sending
When available, a few participants also made use of screenshot prevention in which screenshots result in a blank screen or deterrence messaging from a platform. P15 described using a platform because, “you couldn’t also take like screenshot, like it wasn’t allowed like like how it is [on a banking app]." Since this feature was not readily available in most platforms used for sharing content (see Appendix A), it was also commonly mentioned by others as a feature they wished more platforms would implement it as a means of increasing friction, despite many noting that it could be bypassed (e.g. capturing content through another device).
@@ 8.1.3 After Sharing
Participants proactively used strategies to prevent being identified in the event of NDII. This may be to create plausible deniability (e.g., against blackmail) and to protect against additional harm (e.g., doxxing, harassment). In many cases, participants employed these strategies with the understanding that once content was shared, it was out of their control. “I know that I don’t have full control over where these images will go, if people can trace them back to me" (P26).
@@
Metadata Removal. Six participants used metadata removal to eliminate extra information that is added when an image is created (e.g., location, date/time of creation). However, they generally expressed that they would prefer for it to be automatically done by a platform: “[it] almost feels like that should be the default when you share any image anyway” (P6).
8.3 Defending Against Non-Recipients
@@
Consent. Many participants mentioned the importance of consent prior to sharing intimate content. A few specifically mentioned asking for consent before sharing intimate content in order to check if a recipient was alone or in a private area to prevent shoulder surfing, “I’ll always ask first and wait for their response before I send something because you never know what people are doing” (P29).
Masking Content. Another strategy a few participants used to protect themselves against shoulder surfers was to mask it upon sending it. For the two participants who used this feature, a recipient would receive a blurred image that they would then have to click on to reveal the contents. This would signal that the content was sensitive and should be viewed in a private location. Though this feature was not commonly mentioned, it is also not readily available in most platforms (see Section 9.1).
End-to-End Encryption. End-to-end encryption (E2EE) ensures that only a sender and their intended recipient(s) can view a message. Once an E2EE message has been delivered, it carries similar risks to unencrypted messages: the recipient can still store or share message contents (some platforms may auto-download images onto devices). Over a quarter of participants actively used E2EE or considered it to be useful (excluding those who happened to use platforms with E2EE but did not mention it). For a few, E2EE alleviated their concerns about unwanted access by a platform or government entity: “There are basically two possible attackers. Like either the person I’m sending this to or someone from the outside trying to break into systems… I mostly care about someone from the outside able to steal pictures” (P26). P10 added that they valued E2EE because, “if there’s anything I don’t want, it’s the government knowing details of my personal life and seeing my tits.” On the other hand, a sixth of participants did not trust E2EE, believing that it could be broken if needed: “If someone wants the data, they will get their hands on it anyway. And if that someone is the police, they will just force companies to give that to them. No matter how well encrypted it is” (P50).
@@ 8.4 Comparing Participant Groups
We encountered narratives in which experiencing harm had increased fears of NDII: that, "I’m just really paranoid… and there’s nothing I can really do about it… Especially after it’s already happened" (P35). We also observed that some victim-survivors implemented additional strategies after experiencing harm. As P21 explained, “What happened to me was forever… And so I was more knowledgeable about what was going on, about the privacy of these pictures.” P43 shared, "future photos… were really less revealing than they used to… you cannot recognise me." Otherwise, we did not observe notable differences in practices. Since we interviewed victim-survivors who continued sharing intimate content, many expressed that they resolved not to let shame and stigma stop them (and for some sex workers, they were unable to due to financial reasons). For example, P18 shared:
"I’m having fun. I’m an adult… I don’t want to allow shame or stigma to prevent me from living my life and exploring my sexuality."
When comparing commercial and recreational participants, commercial participants shared content in all the same contexts along with specific platforms to sell content and/or advertise services. However, sex workers were more constrained by the preferences of their clients when selecting a sharing method. Commercial participants may also prefer sex work specific platforms since they may otherwise perceive or experience hostility. We also observed that sex workers were less worried about unintentional discovery, either since the people around them already knew they were sex workers or because they employed more strategies to prevent this.
Generally, we observed that while recreational participants adopted many strategies, sex-working participants adopted more (in particular, watermarking, removing metadata, and using safe storage methods). They were also more likely to set explicit boundaries around sharing or storing. This greater protective action may be due to both differences in threat models (e.g. greater personal/physical risk, stigma) as well as better advice and educational support (see Section 9.2).